OpenStudy (anonymous):

Easy Medal!! Simplify 6 to the negative 3rd power over 6 to the 5th.

OpenStudy (anonymous):

@CUTEYPIE137

OpenStudy (anonymous):

Like this? \[\frac{ 6^{-3} }{ 6^5 }\]

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

\[\frac{ \frac{ 1 }{ 216 } }{ 7776 } = \frac{ 1 }{ 1679616 }\]

OpenStudy (anonymous):

Simplify 6 to the negative 3rd power over 6 to the 5th. 68 6−2 1 over 6 to the 8th power 1 over 6 to the 2nd power

OpenStudy (anonymous):

@saseal which one of those would it be

OpenStudy (anonymous):

whats the question?

OpenStudy (anonymous):

those are the answer choices to the question I asksed for help with :)

OpenStudy (anonymous):

\[6^{8}\] 6 to the 8th power \[6^{-2}\] 1 over 6 to the 2nd power so its 2nd

OpenStudy (anonymous):

its the second answer? @saseal

OpenStudy (anonymous):

yes